### v1

# DREAMS: A python framework to train deep learning models with model card reporting for medical and health applications

Abstract
Electroencephalography (EEG) data provides a non-invasive method for researchers and clinicians to observe brain activity in real time. The integration of deep learning techniques with EEG data has significantly improved the ability to identify meaningful patterns, leading to valuable insights for both clinical and research purposes. However, most of the frameworks so far, designed for EEG data analysis, are either too focused on pre-processing or in deep learning methods per, making their use for both clinician and developer communities problematic. Moreover, critical issues such as ethical considerations, biases, uncertainties, and the limitations inherent in AI models for EEG data analysis are frequently overlooked, posing challenges to the responsible implementation of these technologies. In this paper, we introduce a comprehensive deep learning framework tailored for EEG data processing, model training and report generation. While constructed in way to be adapted and developed further by AI developers, it enables to report, through model cards, the outcome and specific information of use for both developers and clinicians. In this way, we discuss how this framework can, in the future, provide clinical researchers and developers with the tools needed to create transparent and accountable AI models for EEG data analysis and diagnosis.
keywords:
Electroencephalography , model card , deep learning , Python
journal: SoftwareX
Current code version v0.0.3
Permanent link to code/repository used for this code version https://github.com/LucidJun/DREAM
Legal Code License Released under BSD 3-Clause License
Code versioning system used git
Software code languages, tools, and services used Python, Java, CSS, SCSS
If available Link to developer documentation/manual https://dreams-mc.readthedocs.io/en/latest/index.html
Support email for questions rabindra@oslomet.no
Table 1: Code metadata of the Python library DREAMS.
1 Motivation and significance
Existing frameworks dedicated to analyzing EEG data are relatively scarce and those that do exist tend to either cover a wide range of topics inadequately or prioritize either the initial steps of data preparation or the utilization of advanced machine learning techniques [1]. Preprocessing-centric frameworks offer tools tailored to tasks such as noise removal, artifact correction and data segmentation, but may lack advanced analysis capabilities [2]. Deep learning-centric frameworks emphasize using sophisticated algorithms for tasks like classification and feature extraction, assuming preprocessed data [3]. Indeed, this lack of comprehensive frameworks presents challenges for researchers and clinicians seeking effective tools for EEG data analysis.
Furthermore, in the realm of EEG data analysis, ethical considerations play a critical role [4], requiring strict adherence to ethical standards throughout the process of data collection and analysis. This includes securing informed consent from participants and safeguarding their privacy. Biases [5] represent a crucial concern in EEG data analysis, as they can distort results if left unaddressed. These biases, which include selection bias and interpretation bias, must be carefully managed to ensure accurate findings. Uncertainties arising from the complexity of EEG signals present another significant factor that requires acknowledgment and quantification to ensure reliable interpretation. AI models utilized in EEG analysis have inherent limitations, such as interpretability issues, which must be accounted for to prevent misleading conclusions and ensure responsible research and clinical practices [6]. Unfortunately, it is all too common for these critical considerations to be disregarded or given insufficient attention.
Without proper documentation, it becomes difficult to evaluate reliability of the models, effectiveness and potential biases, which are critical factors in ensuring safe and ethical use.
To address the challenges above, we introduce Deep REport for AI ModelS (DREAMS), a Python framework for documenting the performance characteristics of trained machine learning (ML) and deep learning (DL) models. Table 1 provides the main metadata of the code of DREAMS.
2 Related work
Some previous works must be highlighted, which will better contextualize and motivate the novelty of DREMS. Moons et al. [7] introduced the Transparent Reporting of Multivariable Prediction Model for Individual Prognosis or Diagnosis (TRIPOD) statement, which aimed to enhance the reporting of prediction models. The TRIPOD statement is a checklist comprising items essential for the transparent reporting of prediction model studies. This explanation document details the rationale behind each item, clarifies their meanings, and serves as a valuable reference for considerations when reporting the design and analysis of prediction model studies.
Sendak et al. [8] introduced the Model Facts label, designed for clinicians who make decisions supported by machine learning models. Its purpose is to collect relevant, actionable information into a single page. Its major sections include the model name, locale and version, a summary of the model, the mechanism of risk score calculation, validation and performance metrics, usage directions, warnings and additional information.
Crisan et al. [9] explored interactive model cards, which enhance traditional static model cards by allowing users to explore documentation and interact with models. It includes a conceptual study with ML, NLP and AI Ethics experts, followed by an evaluative study with non-expert analysts.
Mitchell et al. [10] presented a framework for transparent model reporting, primarily targeting human-centered models in computer vision and natural language processing, though it can be applied to any machine learning model. They illustrated the concept using model cards for two examples: one that detects smiling faces in images and another that identifies toxic comments in text. All of the architectures mentioned have certain limitations: they may oversimplify complex details and are often tailored to specific settings. Some of these approaches require technical infrastructure for implementation and others are presented as guidelines on particular domains, limiting their broader applicability.
Figure 1: Overview of DREAMS overall architecture.
3 Software description
To illustrate the concept of DREAMS, we present a case study in which model cards (MC) are utilized to highlight the key aspects of the data, model, and performance metrics, as well as the limitations encountered in the study. MCs serve as a detailed documentation tool, emphasizing the characteristics and context of the data used, the specific features and architecture of the model and reports its performance. Moreover, it can include information about limitations or challenges identified during the research. The use of MCs when developing DL underscores the importance of transparency and accountability in model development and deployment.
3.1 Software architecture
DREAMS, developed as a Python package, offers a comprehensive framework aimed at simplifying the process of creating comprehensive model cards for deep learning models. Model cards (MC) serve as vital documentation tools, outlining the performance metrics, usage guidelines and limitations of AI models. They play a pivotal role in enhancing transparency and promoting the ethical utilization of AI technologies.
The DREAMS architecture is modular, comprising several key components. It starts with exploratory data analysis (EDA) on the provided dataset, with the visualizations saved in a designated folder. As the model progresses through training and validation, this folder is further enriched with additional plots, such as loss, accuracy and confusion matrix. Finally, the model card function is invoked, using a configuration file that contains all the necessary details to generate the model card. An overview of the architecture of DREAMS’ library is shown in Figure 1.
3.2 Software functionalities
3.2.1 Data preprocessing
The pipeline typically begins with data collection and preprocessing, where the raw data is gathered from various sources and prepared for model input. Preprocessing steps may include normalization, data augmentation and transformation into formats that are compatible with the model, such as converting images to tensors or encoding text data. Then, the data is usually divided into training and validation sets. The training set is used to train the model, while the validation set assists fine-tune the hyperparameters of the model. To ensure that every step in this process is well-documented and reproducible, we systematically store all critical details about the data, including information on preprocessing, in a designated folder. In a YAML file, we define the exact path to this folder, ensuring that the information is easily accessible and organized for future reference or use. This DREAMS framework setup not only facilitates transparency but also enables efficient tracking of data-related processes throughout the entire model development pipeline.
3.2.2 Model training and evaluation
The training pipeline of a deep learning model is a comprehensive process that transforms raw data into a well-trained model capable of making accurate predictions or classifications. During the training phase, the model processes batches of data iteratively, adjusting its internal parameters to minimize the error between its predictions and the actual outcomes. This is achieved through an optimization algorithm like stochastic gradient descent, which updates the model’s parameters based on the gradients of the loss function. Throughout the training pipeline, performance metrics such as accuracy, loss, precision, recall, and F1-score are monitored to assess the model’s progress. Visualization tools such as confusion matrices are often used to gain deeper insights into the model’s performance. The model is then ready for inference, where it can be used to make predictions on new, unseen data.
At this stage, we meticulously document all the paths to vital information related to both training and evaluation within the YAML file. This DREAMS comprehensive record encompasses not only paths to graphics but also to other essential data, ensuring that every critical component is easily accessible and well-organized. This approach ensures efficient management and retrieval of information throughout the training and evaluation processes.
3.2.3 Model card generation
Once all the necessary information for generating the model card has been gathered, we proceed to invoke the generate_modelcard() function. This function requires three input parameters: the path to the YAML configuration file, which contains the essential details for the model card; the output path, where the generated model card will be saved; and the version number of the model card, which helps track its iterations and updates.
An example of a snippet code using generate_modelcard() function is as follows:
1# model_card library
2from dreams_mc.make_model_card import generate_modelcard
3
4config_file_path="./config.yaml"
5output_path= "./logs/model_card.html"
6version_num = ’1.0’
7
8# Generate the model card
9generate_modelcard(config_file_path, output_path, version_num)
4 An illustrative example for EEG data analysis
To exemplify our approach, we utilize the training pipeline of a deep learning model in conjunction with the Finer-grained Affective Computing EEG Dataset (FACED) dataset [11], which is compatible with TorchEEG  [^1] https://torcheeg.readthedocs.io/en/latest/.
The EEG data were collected from 123 subjects using 32 electrodes positioned according to the international 10–20 system. During the recording sessions, participants viewed 28 different video clips designed to elicit various emotional responses. These video clips spanned nine distinct emotion categories: amusement, inspiration, joy, tenderness, anger, fear, disgust, sadness, and neutral emotion.
The model described in this example classifies EEG data into three main categories: negative (0), which includes emotions such as anger, disgust, fear, and sadness; neutral (1); and positive (2), which includes emotions like amusement, inspiration, joy and tenderness.
To prepare the data for this classification, it is processed through online transformation into tensors and a 2222D format. Then, the labels are processed by adjusting their values accordingly. Following this preprocessing, the dataset is divided into training and validation sets. Subsequently, DataLoaders are created for each set, facilitating the management and batching of data for the training and validation processes. Implementing these steps can be as follows:
1data_folder= "./processed_data_Face/Processed_data"
2
3dataset = FACEDDataset(root_path= data_folder,
4 online_transform=transforms.Compose(
5 [transforms.ToTensor(),
6 transforms.To2d()]),
7 label_transform=transforms.Compose([
8 transforms.Select(’valence’),
9 transforms.Lambda(lambda x: x + 1)
10 ]))
11...
12
13# Dataset Spliting
14train_dataset, val_dataset = random_split(dataset, [num_train_samples, num_val_samples])
15
16# DataLoaders creation
17train_loader = DataLoader(train_dataset, batch_size=batch_size, shuffle=True)
18val_loader = DataLoader(val_dataset, batch_size=batch_size, shuffle=False)
Next, we set up the TSception model [12] and define its hyperparameters:
1#Initiate the model
2model = TSCeption(num_classes=9,
3 num_electrodes=30,
4 sampling_rate=250,
5 num_T=15,
6 num_S=15,
7 hid_channels=32,
8 dropout=0.5)
TSception is a multi-scale convolutional neural network designed to classify emotions from EEG data. It incorporates dynamic temporal layers, asymmetric spatial layers and high-level fusion layers, enabling the model to learn discriminative representations across both time and channel dimensions simultaneously.
Following this, we proceed to configure the training and validation processes utilizing PyTorch. This involves setting up the necessary parameters and components for training the model, such as defining the loss functions, optimizers, and evaluation metrics. We establish the procedures for validating the model’s performance, including managing the data flow, monitoring metrics, and adjusting the training process as needed to ensure effective learning and evaluation. A standard implement for the training and validation configuration is:
1# Model Training
2valid_acc_max = 0.0
3trained_model, accuracy_stats, loss_stats = train(220, valid_acc_max, model, optimizer, criterion, scheduler, train_loader, val_loader, "./logs/current_checkpoint.pt", "./logs/best_model.pt", start_epoch=1)
4
5# plotting and saving training and validation plots
6plot_training_validation_stats(accuracy_stats, loss_stats, save_dir=’./logs’)
7
8# Load the best model for evaluation
9checkpoint_path = ’./logs/best_model.pt’
10checkpoint = torch.load(checkpoint_path, map_location=torch.device(’cpu’))
11model.load_state_dict(checkpoint[’state_dict’])
Upon completion of these procedures, we utilize a custom plotting function to create and store plots depicting accuracy, loss and the confusion matrix in the specified directory:
1plot_metrics_table(results, model_name=’TSception’,save_path="./logs/table.png")
2plot_confusion_matrix(results[’confusion_matrix’], class_names,save_path="./logs/cm.png")
3plot_confidence_intervals(precision=results[’precision’], recall=results[’recall’],
4 accuracy=results[’accuracy’], f1=results[’f1_score’],
5 n=len(val_loader.dataset), save_path="./logs/CI_plot.png",
6 confidence=0.95)
Lastly, we call the generate_model_card() function, providing it with the path to the YAML configuration file  [^2] https://github.com/LucidJun/DREAMS/tree/main/template, the output path for saving the model card, and the version number of the model card:
1# Generating model card
2print("GeneratingModelCard...")
3config_file_path = ’./config.yaml’
4output_path = ’./logs/model_card.html’
5version_num = ’1.0’
6generate_modelcard(config_file_path, output_path, version_num)
Figure 2 provides a detailed overview of the generated model card for this example. This model card encapsulates the key information and metrics resulting from the model’s training and evaluation processes, offering a comprehensive summary of the model’s performance, data characteristics and other relevant details.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Model cards generated for the example described in Sec. 4.
5 Impact
In the context of EEG applications, DREAMS offers a smooth solution for generating comprehensive model cards, providing detailed sections that include overviews of EEG analysis projects, dataset descriptions specific to EEG recordings, results with corresponding plots (such as EEG waveforms or brain activity patterns), uncertainty estimations related to signal interpretation, and relevant references. Additionally, users can customize the content of these model cards using a YAML configuration file, enabling adaptability to various EEG-based projects and models, ensuring flexibility in presenting key insights and findings.
DREAMS is entirely model-agnostic, deliberately engineered to be compatible with any type of model. Its design ensures versatility, allowing it to generate model cards regardless of the underlying model architecture or type. This flexibility is a fundamental aspect of DREAMS, enabling it to adapt to various machine learning and deep learning models, including but not limited to convolutional neural networks (CNNs), recurrent neural networks (RNNs), transformer models and random forests. Whether it’s a classification, regression or any other type of model, DREAMS generates comprehensive model cards, providing valuable insights and documentation irrespective of the model’s complexity or structure. The model-agnostic nature of DREAMS enables users to apply it across a diverse array of projects and applications, ensuring its relevance and usefulness in various machine-learning endeavors.
This versatility allows DREAMS to find applicability across a wide range of domains where transparency and ethical use of AI are paramount. For instance, within the healthcare sector [13, 14], DREAMS could be utilized to document the performance and limitations of deep learning models used for medical analysis. Through the generation of detailed model cards, healthcare practitioners and regulatory bodies can gain insights into the model’s accuracy, potential biases, and areas of uncertainty, facilitating informed decision-making and ensuring patient safety.
Similarly, in the financial sector [15, 16], DREAMS can aid in documenting AI models used for credit risk assessment or fraud detection, providing transparency into the model’s decision-making process and potential biases. This transparency is essential for ensuring fairness and accountability in algorithmic decision-making processes.
Furthermore, in intelligent transportation systems [17] such as, autonomous vehicles, DREAMS can be employed to document the performance and safety considerations of deep learning models used for object detection and navigation, promoting trust and transparency among regulators and the general public.
DREAMS can be effortlessly installed and utilized via PyPI, ensuring convenient access for users. The project’s code has been shared on GitHub, allowing for collaborative contributions and enhancements from the community. As an open-source initiative, users can extend and customize the software to meet their specific requirements. Comprehensive documentation is available in English to assist users in understanding and utilizing the software effectively. Note that, the maintenance and upkeep of the software are managed by one of the authors, ensuring ongoing support and reliability.
In essence, DREAMS enables the creation of model cards that hold immense value for AI practitioners. These cards provide vital documentation that enhances transparency, facilitates user understanding, and supports responsible deployment by detailing the operational parameters of the model.
6 Conclusions and future perspectives
This paper introduces DREAMS, a Python library designed to simplify the creation of comprehensive model cards for deep learning models.
In fact, DREAMS offers efficient generation of detailed model cards with customizable content through YAML configuration, covering project overviews, dataset descriptions, results with plots, uncertainty estimations, and references. It also supports various plot types, enhancing clarity and comprehension.
As such, DREAMS serves as a valuable tool for promoting transparency and ethical use of AI across various domains, facilitating informed decision-making, and fostering trust in AI systems. Whether in healthcare, finance, intelligent transportation systems, or any other domain where AI is utilized, DREAMS enables the documentation of model performance, usage, and limitations through comprehensive model cards. This documentation not only enhances understanding but also promotes accountability and ensures that AI systems are deployed responsibly and ethically.
Acknowledgements
The authors thank Harry Hallock and Aleksandar Babic (DNV, Norway) for useful discussions and suggestions around the topic of model cards and data cards. This work has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 964220. We conducted experiments on the Experimental Infrastructure for Exploration of Exascale Computing (eX3) system, financially supported by RCN under contract 270053.